solve for x in the following "linear" equation showing all work. do not use graphs to define x. (x/(x^2-1))-((x+3)/(x^2-x))=(-3/(x^2+x))
1
Expert's answer
2012-09-18T10:59:58-0400
(x/(x+1)(x-1)) – ((x+3)/x(x-1))+3/x(x+1)=0
(x2-(x+3)(x+1)+3(x-1))/(x(x-1)(x+1))=0
X=-6 , X≠0, X≠1, X≠-1